# Pre-trained Encoder Inference: Revealing Upstream Encoders In Downstream Machine Learning Services

### v1

@@ Abstract
Though pre-trained encoders can be easily accessed online to build downstream machine learning (ML) services quickly, various attacks have been designed to compromise the security and privacy of these encoders. While most attacks target encoders on the upstream side, it remains unknown how an encoder could be threatened when deployed in a downstream ML service. This paper unveils a new vulnerability: the Pre-trained Encoder Inference (PEI) attack, which posts privacy threats toward encoders hidden behind downstream ML services. By only providing API accesses to a targeted downstream service and a set of candidate encoders, the PEI attack can infer which encoder is secretly used by the targeted service based on candidate ones. We evaluate the attack performance of PEI against real-world encoders on three downstream tasks: image classification, text classification, and text-to-image generation. Experiments show that the PEI attack succeeds in revealing the hidden encoder in most cases and seldom makes mistakes even when the hidden encoder is not in the candidate set. We also conducted a case study on one of the most recent vision-language models, LLaVA, to illustrate that the PEI attack is useful in assisting other ML attacks such as adversarial attacks. The code is available at https://github.com/fshp971/encoder-inference.
I Introduction
@@
Recent advances of self-supervised learning (SSL) [1, 2, 3, 4, 5, 6] and the availability of large amounts of public unlabeled data have enabled the pre-training of powerful representation encoders. These pre-trained encoders are usually first built by upstream suppliers who control sufficient computational resources and then delivered through online model repositories (e.g., Hugging Face [^1] Hugging Face model repository is at https://huggingface.co/models.) or Encoder-as-a-Service (EaaS) to downstream suppliers to help them quickly build their own machine learning (ML) services. Specifically, one can simply use an encoder to encode downstream training data to informative embeddings and train downstream models upon them with relatively little cost. These downstream models are eventually provided as service APIs to end users.
Figure 1: An illustration of the threat model of the PEI attack. In this model, the adversary has API accesses to a set $E$ of candidate encoders and a targeted downstream ML service built upon a hidden encoder $f^{*}$. The attack goal is to infer whether the hidden encoder $f^{*}$ is in the candidate set $E$, and if so, further infer which candidate $f_{i}\in E$ is the hidden $f^{*}$.
@@
Despite the success of this upstream pre-training, downstream quick-building paradigm, many attacks have also emerged to threaten the security and privacy of pre-trained encoders. Current attacks toward pre-trained encoders can be roughly divided into two categories: The first category aims to attack encoders in the SSL stage via poisoning [7, 8] or backdooring [9, 10, 11, 12] to manipulate their behaviors, while the second one targets trained encoders and aims to extract private information of their training data [13, 14, 15] or unauthorizedly duplicate their encoding functionalities [16, 17, 18]. To mitigate these attacks, various defenses are designed to enhance the robustness of encoders against adversaries [19, 20, 21] or perform source tracing when an attack has occurred [22, 23, 24].
While attacks and defenses toward bare pre-trained encoders on the upstream side have been vastly studied, encoders that have already been deployed in downstream services do not receive equal attention, which may be due to the common belief that they are usually much safer on the downstream side. Concretely, when a pre-trained encoder is deployed in a downstream service, the end users can only see and access the service APIs powered by the downstream model. In this way, the downstream model seems like a shield that isolates the encoder from end adversaries, leaving them unable to interact directly with it using malicious queries. Though such a “downstream protection” is supported by the data processing inequality to some extent, it is still not quantified how much the information of the hidden encoder is sanitized through the downstream model and APIs. Thus, it remains open whether and how a pre-trained encoder deployed in a downstream ML service could be threatened by malicious end users.
Our works. This paper shows that pre-trained encoders can indeed be threatened from the downstream side by introducing a new class of privacy attacks named Pre-trained Encoder Inference (PEI) attack. As illustrated in Figure 1, by providing a set $E$ consisting of several candidate pre-trained encoders, for a targeted downstream ML service that is built upon a hidden encoder $f^{*}$ (note that $f^{*}$ may not in $E$), the PEI attack will infer whether the hidden encoder $f^{*}$ comes from the candidate set $E$. Furthermore, if the hidden encoder is inferred to be from the candidate set, the PEI attack will reveal which candidate encoder is the targeted hidden encoder. A key feature makes our PEI attack very practical in the real world: the adversary only requires API access to candidate encoders and the targeted ML service to perform the PEI attack.
@@
The design of the PEI attack is motivated by the observation that for any certain encoder and a pre-defined embedding, there may exist samples that look different from each other but enjoy embeddings similar to that pre-defined embedding only under such a certain encoder. Therefore, we propose a general framework to implement the PEI attack by first synthesizing a series of PEI attack samples via minimizing the difference of their embeddings and the pre-defined embedding under the certain encoder from the candidate set, and then inferring whether this certain candidate encoder is used by a targeted downstream ML service by evaluating whether these synthesized PEI attack samples can make the targeted service produce a specified behavior determined by that pre-defined embedding. Since in our threat model, the adversary only has API access to candidate encoders, the PEI attack further exploits zeroth-order optimization [25, 26, 27] to synthesize PEI attack samples for different candidate encoders. Besides, it is worth noting that while the synthesis of PEI attack samples depends on specific data modalities, our attack framework makes no other assumptions about the downstream tasks. Thus, in other words, the PEI attack is downstream task-agnostic as long as the input data modality is given, which is another advantage of our attack: once the PEI attack samples are synthesized, the adversary can use them to attack any kind of downstream ML services.
While the aforementioned general attack framework can be applied to encoders on any data modalities, in this work we focus on analyzing two families of encoders: image encoders and text encoders. We conduct experiments to evaluate the PEI attack against different image/text encoders in three downstream tasks: image classification, text classification, and text-to-image generation. Results show that the PEI attack successfully reveals correct hidden encoders from $24$ out of $24$ image classification services, $14$ out of $24$ text classification services, and $4$ out of $4$ text-to-image services. The PEI attack also demonstrates a strong capability in avoiding false-positive predictions: it seldom infers incorrect encoders as the hidden ones even when the correct hidden encoder is not in the candidate set. This feature is handy for adversaries as they do not need to waste time in tackling false-positive signals and thus will always have a good chance to find a victim in a short time after many unsuccessful but quick tries. In addition to the significant attack performance and practicality, the cost of black-box synthesizing PEI attack samples is also low, with an estimated price of no more than \$$\$400$ per candidate encoder.
We further argue that the PEI attack is useful in practice by conducting a case study on how the PEI attack can assist adversarial attacks against LLaVA [28, 29], a recent multimodal model. The overall attack goal is to synthesize visually benign adversarial images to induce LLaVA to generate malicious textual content. Concretely, the adversary will first use the PEI attack to reveal the image encoder hidden in LLaVA. With direct access to this hidden encoder, the adversary will then synthesize adversarial images that contain only benign mosaics but enjoy embeddings very similar to images that contain pre-defined harmful texts. As a result, when feeding these synthesized mosaic images to LLaVA, the model will generate texts that are similar to the pre-defined harmful contents. Such a PEI-assisted adversarial attack can be used to stealthily spread harmful information such as false medical/health information or hate speeches, since the spreading mediums, i.e., the mosaic adversarial images, do not contain visually harmful information and thus are difficult to be detected and filtered. As a result, there is indeed a practical need to develop general defenses against the PEI attack.
Finally, as a preliminary investigation, we discuss potential defenses against the PEI attack, which include methods that sanitizing or detecting PEI attack samples, or re-designing downstream models to be natively robust to the PEI attack.
Contributions. In summary, our work makes the following contributions: (1) We unveil a new vulnerability of pre-trained encoders named Pre-trained Encoder Inference (PEI) attack that can reveal what encoder is secretly used by a targeted downstream ML service (Section III-B). (2) We propose a general framework to implement the PEI attack in a black-box and downstream task-agnostic manner (Section IV), and instantiate it for image encoders and text encoders (Section V). (3) We empirically verify the effectiveness of the PEI attack on three families of downstream tasks, which are image/text classifications (Section VI), and text-to-image generation (Section VII). (4) We conduct a case study on LLaVA to show how the PEI attack can assist adversarial attacks against multimodal models (Section VIII). The result demonstrates the usefulness of the PEI attack and thus calls for a practical need to develop defenses against the PEI attack. (5) Finally, we discuss several potential defenses against the PEI attack (Section IX).
II Related Works
Privacy attacks against pre-trained encoders. For data privacy attacks, many efforts have been made to design MIA against encoders. When the pre-training strategy is known, Liu et al. [13] successfully performs membership inference attacks (MIA) against targeted encoder based on membership scores calculated with the pre-training loss function and shadow training techniques. He et al. [14] extends MIA against semi-supervisedly learned encoders. Zhu et al. [15] studies a more challenging setting where the adversary has no knowledge about pre-training algorithms and proposes to calculate the embedding similarity between global and local features as the membership score for each sample. Wen et al. [11] and Feng and Tramèr [12] show that backdoored encoders can increase the privacy vulnerabilities of downstream models built upon them. Besides, for model privacy attacks, studies mainly focus on stealing the powerful encoding functionality of encoders. By simply querying the targeted encoder and collecting return embeddings, Liu et al. [16] succeeds in retraining an encoder that reproduces the functionality of the target with a reasonable query budget. Sha et al. [17] adopted a contrastive learning-based method to further improve the encoder stealing performance. However, all these data/model privacy attacks usually require detailed information about the downstream tasks, such as the number of classes in downstream classification tasks or prior knowledge of potential targeted training data, which may limit their practicality in the real world. As a comparison, our PEI attack can work in a downstream task-agnostic manner.
III Preliminaries
III-A Pre-trained Encoders
@@
A pre-trained encoder is a function $f:\mathcal{X}\rightarrow\mathcal{E}$ that can encode any sample point $x\in\mathcal{X}$ to an informative embedding vector $f(x)$. This powerful encoding ability can be used to facilitate building downstream ML services, (e.g., image/text classification, and text-to-image generalization). Specifically, suppose there is a downstream task-specific dataset $D=\{(x_{i},y_{i})\}_{i=1}^{|D|}$, where $x_{i}\in\mathcal{X}$ is the $i$-th downstream sample and $y_{i}\in\mathcal{Y}$ is its corresponding label. To build an ML service to handle the downstream task, the service supplier will encode each downstream sample $x_{i}$ to $f(x_{i})$, and train a downstream model $h_{\theta}$ on these encoded embeddings as follows,
@@
where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a downstream task-dependent loss function and $\theta$ is the downstream model parameter. After that, the downstream ML service will be made online and provide API access to end users, in which the API is powered by the composite function $h_{\theta}(f(\cdot)):\mathcal{X}\rightarrow\mathcal{Y}$. Whenever a query $x$ comes, the service API will return $h_{\theta}(f(x))$ as the response.
III-B Threat Model
The threat model of the PEI attack is illustrated in Figure 1, which consists of two parties: (1) a targeted downstream ML service $g:\mathcal{X}\rightarrow\mathcal{Y}$, and (2) the adversary. We next introduce them in detail.
Downstream ML service $g$. The service can be seen as a composite function $g(\cdot):=h_{\theta}(f^{*}(\cdot))$, where $f^{*}$ is the encoder secretly used in the downstream service and $h_{\theta}$ is the downstream model trained following the pipeline described in Section III-A. It is notable that this targeted service $g$ can only be accessed through APIs: for any query sample $x$ from end-users, the service will only return $g(x)$ as the response. Other operations are not allowed.
Adversary’s goal. The overall goal of the adversary is to reveal the encoder $f^{*}$ hidden by the targeted service $g$. Concretely, suppose the adversary holds a set $E=\{f_{1},\cdots,f_{N}\}$ consisting of $N$ publicly accessible pre-trained encoders that may come from online model repositories or EaaSs. Then, the adversary aims to: (1) determine whether the hidden encoder $f^{*}$ comes from the candidates set $E$, and (2) if it is determined that $f^{*}\in E$, infer which candidate $f_{i}\in E$ is the hidden $f^{*}$.
@@ III-B Threat Model
  • API access to the targeted service. The adversary can query the targeted service $g$ with any sample $x$ and receive the return $g(x)$. It is notable that the adversary has no prior knowledge about the downstream model $h_{\theta}$ and the hidden encoder $f^{*}$ and is not allowed to directly interact with the hidden $f^{*}$ through $g$.
  • API access to candidate encoders. For each candidate $f_{i}\in E$, the adversary can query it with any sample $x$ and receive the return $f_{i}(x)$. However, the adversary could not access the model parameter of $f_{i}$. We argue that the size of the candidate set would not be too large, as real-world applications usually tend to use EaaSs/pre-trained encoders provided by a few reputable tech companies (e.g., OpenAI, Hugging Face, Microsoft, and Google).
IV PEI Attack Framework
Figure 2: Examples of the PEI attack against image classification service (ResNet-50 (Hugging Face) + CIFAR-10), text classification service (BERT (base) + TREC), and text-to-image generation service (StableDiffusion v2-1 with OpenCLIP ViT-H/14). The high-level attack idea is to first synthesize PEI attack samples that enjoy embeddings similar to that of a pre-defined objective sample, and then perform the inference attack by evaluating whether the synthesized attack samples can make downstream targeted service produce behaviors similar to that of the objective sample.
In this section, we present a general framework to perform the PEI attack. At a high level, our designed framework is motivated by the hypothesis that for a certain encoder and a pre-defined embedding, there may exist samples that look different but enjoy embeddings similar to that pre-defined one only under this certain encoder. Further, when the encoder is changed, the embeddings of these samples will again become different. We named such kind of samples as PEI attack samples corresponding to the certain encoder, and they will be the key ingredient of our attack framework.
The idea is that for a targeted downstream service that is built upon this certain encoder, the corresponding PEI attack samples are very likely to make the downstream model produce a specific behavior determined by the pre-defined embedding. Besides, when the hidden encoder behind the targeted service is not the aforementioned certain encoder, such a specific behavior is less likely to be produced by the downstream model. Therefore, the adversary can exploit this behavior discrepancy to perform the PEI attack against downstream ML services. Such an idea is also illustrated in Figure 2, which includes examples of the PEI attack against image classification, text classification, and text-to-image generation.
@@ IV PEI Attack Framework
3:// PEI Attack Samples Synthesis Stage
4:for $i$ in $1,\cdots,N$ do
5:     Synthesize $\{x^{(atk)}_{i,j,k}\}_{k=1}^{M_{2}}$ via first randomly initializing them and then minimizing Eq. (1) based on $f_{i}$, $x^{(obj)}_{j}$, and a zeroth-order optimizer.
6:end for
7:// Hidden Encoder Inference Stage
8:for $i$ in $1,\cdots,N$ do
9:     Calculate the PEI score $\zeta_{i}$ following Eq. (2) based on $g$, $f_{i}$, $\{x^{(obj)}_{j}\}_{j=1}^{M_{1}}$ and $\{x^{(atk)}_{i,j,k}:j\leq M_{1},k\leq M_{2}\}$.
10:end for
11:Calculate $z_{1},\cdots,z_{N}$ following Eq. (4).
12:$I\leftarrow\{i:z_{i}>1.7\}$
13:if $|I|=1$ then
14:     return $\hat{f}^{*}\leftarrow f_{i^{*}}$triangleright $i^{*}$ is the only element in set $I$.
15:end if
16:return $\hat{f}^{*}\leftarrow\emptyset$
IV-A General Framework
PEI attack samples synthesis stage. In this stage, the adversary first collects $M_{1}$ samples $x_{1}^{(obj)},\cdots,x_{M_{1}}^{(obj)}$ named objective samples from public data domains. Then, for each encoder-sample pair $(f_{i},x^{(obj)}_{j})$ where $f_{i}\in E\ (1\leq i\leq N)$ is the $i$-th candidate encoder and $x^{(obj)}_{j}\ (1\leq j\leq M_{1})$ is the $j$-th objective sample, the adversary synthesizes a set of $M_{2}$ PEI attack samples $\{x^{(atk)}_{i,j,k}\}_{k=1}^{M_{2}}$. Each PEI attack sample $x^{(atk)}_{i,j,k}$ for the pair $(f_{i},x^{(obj)}_{j})$ is first randomly initialized and then obtained via minimizing the following squared loss $\mathcal{L}_{i,j}$,
Intuitively, Eq. (1) aims to make the embeddings of the PEI attack sample $x^{(atk)}_{i,j,k}$ and the objective sample $x^{(obj)}_{j}$ under the candidate encoder $f_{i}$ similar to each other in $l_{2}$-distance. To minimize Eq. (1), a naive solution is to use gradient descent methods. However, since our threat model only assumes black-box access to the candidate encoder $f_{i}$, the adversary is not able to calculate first-order gradients of Eq. (1), which thus makes simple gradient descent methods become invalid. We skip this for now and will show how to tackle Eq. (1) via zeroth-order optimization [27, 26, 25] in Section V.
Hidden encoder inference stage. In this stage, for the pre-trained encoder $f^{*}$ hidden in the targeted downstream ML service $g(\cdot):=h_{\theta}(f^{*}(\cdot))$, the adversary exploits the synthesized PEI attack samples to infer (1) whether $f^{*}\in E$ or not and (2) if so, which $f_{i}\in E$ is $f^{*}$. We argue that these two goals can be solved simultaneously. Concretely, we propose to first calculate $N$ PEI scores $\zeta_{1},\cdots,\zeta_{N}$ for the $N$ candidates, where each $\zeta_{i}$ is calculated based on objective samples $x^{(obj)}_{1},\cdots,x^{(obj)}_{M_{1}}$ and PEI attack samples $\{x^{(atk)}_{i,j,k}:j\leq M_{1},k\leq M_{2}\}$ corresponding to the candidate $f_{i}\in E$ as follows,
@@ IV-A General Framework
where $\ell_{sim}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a task-dependent function that measures the similarity of behaviors produced by the targeted downstream service $g$. A larger output $\ell_{sim}(\cdot,\cdot)$ means the corresponding two inputs are more similar to each other. Then, among these $N$ PEI scores $\zeta_{1},\cdots,\zeta_{N}$, if there happens to be a single PEI score, denoted as $\zeta_{i^{*}}$, that is significantly higher than others, one can thus conclude that $f^{*}\in E$ and $f^{*}=f_{i^{*}}$, otherwise conclude $f^{*}\notin E$. After that, the aforementioned two PEI attack goals have been accomplished.
Based on the above analysis, the remaining task to complete the PEI attack is to find a way to assess whether a given PEI score is truly significantly higher than others. Here we leverage a simple yet efficient one-tailed $z$-test to realize our goal. Specifically, for each PEI score $\zeta_{i}$, if the candidate encoder $f_{i}$ is not the hidden one $f^{*}$, then the score $\zeta_{i}$ would be relatively low. As a result, we can check whether $\zeta_{i}$ is significantly high by testing the following null hypothesis,
$H^{(i)}_{0}:$ The PEI score $\zeta_{i}$ is NOT calculated based on
PEI attack samples for the hidden encoder $\text{\it PEI attack samples for the hidden encoder $f^{*}$}.$ (3)
If $H^{(i)}_{0}$ is rejected, then $\zeta_{i}$ will be determined to be significantly high and $f_{i}$ will be inferred as the hidden $f^{*}$.
To use the one-tailed $z$-test to evaluate the null hypothesis, one needs to construct the following statistic z-score,
$z_{i}=\frac{\zeta_{i}-\mathrm{E}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}}{\mathrm{SD}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}},$ (4)
where $\mathrm{E}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}$ and $\mathrm{SD}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}$ are the mean and the sample standard deviation of the $N$ candidate PEI scores $\zeta_{1},\cdots,\zeta_{N}$. If the $z_{i}$ is above a chosen threshold, then the null hypothesis $H^{(i)}_{0}$ will be rejected. We further assume that the distribution of PEI scores calculated based on non-hidden encoders’ PEI attack samples can be approximated by a normal distribution. In this case, if we chose $z_{i}>1.7$ as the rejection criterion, then it will lead to a one-sided $p$-value, which is also the false-positive rate, of $4.45\%$. Since such a $p$-value is smaller than the typical statistical significance level threshold (which is $5\%$), we thus set the criterion as $z_{i}>1.7$, which eventually leads to the following inference attack,

### v2

@@ Abstract
Pre-trained encoders available online have been widely adopted to build downstream machine learning (ML) services, but various attacks against these encoders also post security and privacy threats toward such a downstream ML service paradigm. We unveil a new vulnerability: the Pre-trained Encoder Inference (PEI) attack, which can extract sensitive encoder information from a targeted downstream ML service that can then be used to promote other ML attacks against the targeted service. By only providing API accesses to a targeted downstream service and a set of candidate encoders, the PEI attack can successfully infer which encoder is secretly used by the targeted service based on candidate ones. Compared with existing encoder attacks, which mainly target encoders on the upstream side, the PEI attack can compromise encoders even after they have been deployed and hidden in downstream ML services, which makes it a more realistic threat. We empirically verify the effectiveness of the PEI attack on vision encoders. we first conduct PEI attacks against two downstream services (i.e., image classification and multimodal generation), and then show how PEI attacks can facilitate other ML attacks (i.e., model stealing attacks vs. image classification models and adversarial attacks vs. multimodal generative models). Our results call for new security and privacy considerations when deploying encoders in downstream services. The code is available at https://github.com/fshp971/encoder-inference.
1 Introduction
@@
Recent advances of self-supervised learning (SSL) [15, 54, 28, 70, 55, 10] and the availability of large amounts of public unlabeled data have enabled the pre-training of powerful representation encoders. These pre-trained encoders are usually first built by upstream suppliers who control sufficient computational resources and then delivered through online model repositories (e.g., Hugging Face) or Encoder-as-a-Service (EaaS) to downstream suppliers to help them quickly build their own machine learning (ML) services. Specifically, one can simply use an encoder to encode downstream training data to informative embeddings and train downstream models upon them with relatively little cost. These downstream models are eventually provided as service APIs to end users.
@@
Despite the success of this upstream pre-training, downstream quick-building paradigm, many attacks also emerge to threaten pre-trained encoders and further compromise the security and privacy of downstream ML services. These encoder attacks typically begin by compromising encoders in the pre-training stage via poisoning [6] or backdooring [35]. Then, when an attacked encoder is deployed in a downstream ML service, it will enable the adversary to manipulate the behavior of the downstream model [42, 60] or leak private information of downstream data [67, 21]. To mitigate these attacks, efforts have been made during the encoder pre-training stage to enhance the robustness of encoders against adversaries [49, 61] or perform source tracing when an attack has occurred [46, 13, 18].
Figure 1: Illustration of how the PEI attack can threaten downstream ML services. Step 1: Using the PEI attack to reveal the encoder hidden in the targeted downstream service. Step 2: Exploiting the revealed encoder to conduct other ML attacks, e.g., model stealing and adversarial attacks, against the original targeted service.
This paper focuses on a more challenging setting where a clean pre-trained encoder (i.e., not affected by encoder attacks) has already been deployed into the targeted service and the adversary only has API access to this service. Under this setting, pervious poisoning/backdooring-based encoder attacks, which need to access and modify the pipeline of building downstream ML services, all become invalid. In this sense, both encoders and downstream ML services seem to be much safer. Unfortunately, we unveil a new class of encoder privacy attacks named Pre-trained Encoder Inference (PEI) attack, showing that downstream services and their encoders are still vulnerable to adversaries from the downstream side even under such a “safer setting”. As shown in Figure 1, given a targeted downstream ML service, the goal of the new PEI attack is to infer what pre-trained encoder is secretly used by the downstream model. Once the hidden encoder is revealed, the adversary can then exploit public APIs of the hidden encoder to facilitate other ML attacks against the targeted downstream service such as model stealing attacks or adversarial attacks more effectively.
@@
To launch a PEI attack against a targeted downstream ML service that is built upon a hidden encoder $f^{*}$, we assume that the adversary has API accesses to the targeted service and a set $E$ consisting of several candidate pre-trained encoders from online model repositories or EaaSs. The goal of the PEI adversary is to infer: (1) whether $f^{*}\in E$ and (2) which candidate from $E$ is the hidden $f^{*}$. Realizing the PEI attack relies on the observation that for any certain encoder and a pre-defined embedding, there exist samples that look different from each other but enjoy embeddings similar to that pre-defined one only under such a certain encoder. Under this observation, to infer whether a given candidate encoder is used by the targeted service, we propose to first synthesize a series of PEI attack samples via minimizing the difference of their embeddings and the pre-defined embedding under this candidate encoder. Then, we evaluate whether these synthesized PEI attack samples can make the targeted service produce a specified behavior determined by that pre-defined embedding. Notably, our framework makes no assumptions about downstream tasks, and thus can work in a downstream-task agnostic manner as long as the PEI attack samples are synthesized.
We conduct experiments on vision encoders used in two downstream tasks: image classification and multimodal generation (i.e., the LLaVA model [40, 41]), to demonstrate the effectiveness of the proposed PEI attack. For each targeted downstream service, we first conduct the PEI attack against the hidden encoder and demonstrate that the PEI attack successfully reveals hidden encoders in most targeted services, with a cost as low as an estimated \$100 per candidate encoder. Moreover, our attack never makes false-positive predictions, even when the correct hidden encoder is not in the candidate set. Then, we study how the hidden encoder revealed by the PEI attack can facilitate other ML attacks against the targeted service. For the image classification service, we show that the PEI attack can improve model stealing attacks [63] in both accuracy and fidelity by $2\sim 20$ times. For the multimodal generative service, we analyze how the revealed encoder can assist in synthesizing adversarial images to stealthily spread false medical/health information through the targeted service.
Due to space limitations, discussions on related works are included in Appendix A.
2 Preliminaries
2.1 Pre-trained encoders
@@
A pre-trained encoder is a function $f:\mathcal{X}\rightarrow\mathcal{E}$ that can encode any sample $x\in\mathcal{X}$ to an informative embedding vector $f(x)$. This powerful encoding ability can be used to facilitate building downstream ML services. Concretely, suppose there is a downstream task-specific dataset $D=\{(x_{i},y_{i})\}_{i=1}^{|D|}$, where $x_{i}\in\mathcal{X}$ is the $i$-th downstream sample and $y_{i}\in\mathcal{Y}$ is its label. To build an ML service to handle the downstream task, the service supplier will encode each downstream sample $x_{i}$ to $f(x_{i})$, and train a downstream model $h_{\theta}$ on these encoded embeddings as follows,
@@
where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a downstream task-dependent loss function and $\theta$ is the downstream model parameter. After that, the downstream ML service will be made online and provide API access to end users, where the API is powered by the composite function $h_{\theta}(f(\cdot)):\mathcal{X}\rightarrow\mathcal{Y}$. Whenever a query $x$ comes, the API will return $h_{\theta}(f(x))$ as the response.
2.2 Threat model
The threat model of the PEI attack consists of two parties: (1) the targeted downstream ML service $g:\mathcal{X}\rightarrow\mathcal{Y}$, and (2) the PEI adversary.
Targeted downstream ML service $g$. The targeted service can be seen as a composite function $g(\cdot):=h_{\theta}(f^{*}(\cdot))$, where $f^{*}$ is the encoder secretly used in the downstream service and $h_{\theta}$ is the downstream model trained following the pipeline described in Section 2.1. It is notable that this targeted service $g$ can only be accessed through APIs: for any query sample $x$ from end-users, the service will only return $g(x)$ as the response. Other operations are not allowed.
Adversary’s goal. The goal of a PEI adversary is to reveal the encoder $f^{*}$ hidden by the targeted service $g$. Concretely, suppose the adversary holds a set $E=\{f_{1},\cdots,f_{N}\}$ consisting of $N$ publicly accessible pre-trained encoders that may from online model repositories or EaaSs. The adversary aims to: (1) determine whether the hidden $f^{*}$ comes from the candidates set $E$, and (2) if it is determined that $f^{*}\in E$, infer which candidate $f_{i}\in E$ is the hidden $f^{*}$.
@@ 2.2 Threat model
  • API access to the targeted service. The adversary can query the targeted service $g$ with any sample $x$ and receive $g(x)$ as the return. It is notable that the adversary has no prior knowledge about the downstream model $h_{\theta}$ and the hidden encoder $f^{*}$ and is not allowed to directly interact with or modify the hidden encoder $f^{*}$.
  • API access to candidate encoders. For each candidate $f_{i}\in E$, the adversary can query it with any sample $x$ and receive $f_{i}(x)$ as the return. However, the adversary could not access the model parameter of $f_{i}$. We argue that the size of the candidate set does not need be too large, as real-world applications usually tend to use EaaSs/pre-trained encoders provided by a few reputable tech companies (e.g., OpenAI, Hugging Face, Microsoft, and Google).
3 Designing Pre-trained Encoder Inference (PEI) attack
3.1 Intuition of the design
The high-level design of the PEI attack is motivated by the observation that for a certain encoder and a pre-defined embedding, samples always exist that look different but enjoy embeddings similar to that pre-defined one only under this encoder. Further, when the encoder is changed, the embeddings of these samples will again become different. We named such kind of samples as PEI attack samples corresponding to the certain encoder, and they will be the key ingredient of our attack framework.
The idea is that for a targeted downstream service that is built upon this certain encoder, the corresponding PEI attack samples are very likely to make the downstream model produce a specific behavior determined by the pre-defined embedding. Besides, when the hidden encoder behind the targeted service is not the aforementioned certain encoder, such a specific behavior is less likely to be produced by the downstream model. Thus, the adversary can exploit this behavior discrepancy to perform the PEI attack against downstream ML services.
3.2 General attack framework
PEI attack samples synthesis stage. In this stage, the adversary first collects $M_{1}$ samples $\{x_{j}^{(obj)}\}_{j=1}^{M_{1}}$ named objective samples from public data domains. Then, for each encoder-sample pair $(f_{i},x^{(obj)}_{j})$ where $f_{i}\in E$ is the $i$-th candidate encoder and $x^{(obj)}_{j}$ is the $j$-th objective sample, the adversary synthesizes a set of $M_{2}$ PEI attack samples $\{x^{(atk)}_{i,j,k}\}_{k=1}^{M_{2}}$. Each $x^{(atk)}_{i,j,k}$ for the pair $(f_{i},x^{(obj)}_{j})$ is first randomly initialized and then obtained via minimizing the below squared loss $\mathcal{L}_{i,j}$,
Intuitively, Eq. (1) aims to make the embeddings of the PEI attack sample $x^{(atk)}_{i,j,k}$ and the objective sample $x^{(obj)}_{j}$ under the candidate encoder $f_{i}$ similar to each other in $l_{2}$-distance. To minimize Eq. (1), a naive solution is to use gradient descent methods. However, since our threat model only assumes API access to the candidate encoder $f_{i}$, the adversary cannot calculate first-order gradients of Eq. (1), which makes simple gradient descent methods be invalid.
@@ 3.2 General attack framework
3:for $i$ in $1,\cdots,N$ do $\triangleright$ PEI Attack Samples Synthesis Stage Started
4:     Synthesize $\{x^{(atk)}_{i,j,k}\}_{k=1}^{M_{2}}$ via first randomly initializing them and then minimizing Eq. (1) based on $f_{i}$, $x^{(obj)}_{j}$, and a zeroth-order optimizer.
5:end for
6:for $i$ in $1,\cdots,N$ do $\triangleright$ Hidden Encoder Inference Stage Started
7:     Calculate the PEI score $\zeta_{i}$ following Eq. (2) based on $g$, $f_{i}$, $\{x^{(obj)}_{j}\}_{j=1}^{M_{1}}$ and $\{x^{(atk)}_{i,j,k}\}_{j=1,k=1}^{M_{1},M_{2}}$.
8:end for
9:Calculate $z_{1},\cdots,z_{N}$ following Eq. (3).
10:$I\leftarrow\{i:z_{i}>1.7\}$
11:if $|I|=1$ then
12:     return $\hat{f}^{*}\leftarrow f_{i^{*}}$triangleright $i^{*}$ is the only element in set $I$.
13:end if
14:return $\hat{f}^{*}\leftarrow\emptyset$
Hidden encoder inference stage. For the encoder $f^{*}$ hidden in the targeted downstream ML service $g(\cdot):=h_{\theta}(f^{*}(\cdot))$, the adversary exploits synthesized PEI attack samples to infer: (1) whether $f^{*}\in E$ or not and (2) if so, which $f_{i}\in E$ is $f^{*}$. We argue that these two goals can be solved simultaneously. Concretely, we propose to first calculate $N$ PEI scores $\zeta_{1},\cdots,\zeta_{N}$ for the $N$ candidates, where each $\zeta_{i}$ is calculated based on objective samples $x^{(obj)}_{1},\cdots,x^{(obj)}_{M_{1}}$ and PEI attack samples $\{x^{(atk)}_{i,j,k}:j\leq M_{1},k\leq M_{2}\}$ of the candidate $f_{i}\in E$ as follows,
@@
where $\ell_{sim}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a task-dependent function that measures the similarity of behaviors produced by the targeted service $g$. A larger output $\ell_{sim}(\cdot,\cdot)$ means the corresponding two inputs are more similar to each other. Among the $N$ PEI scores, if there happens to be a single score, denoted as $\zeta_{i^{*}}$, that is significantly higher than others, one can thus conclude that $f^{*}\in E$ and $f^{*}=f_{i^{*}}$, otherwise conclude $f^{*}\notin E$. After that, the two PEI attack goals have been accomplished.
Based on the above analysis, the remaining task to complete the PEI attack is to find a way to assess whether a given PEI score is truly significantly higher than others. Here we adopt a simple yet efficient one-tailed $z$-test to realize our goal. Specifically, for each PEI score $\zeta_{i}$, if the candidate encoder $f_{i}$ is not the hidden one $f^{*}$, then the score $\zeta_{i}$ would be relatively low. As a result, we can check whether $\zeta_{i}$ is significantly high by testing the following null hypothesis,
$H^{(i)}_{0}:\ \text{\it The PEI score $\zeta_{i}$ is {NOT} calculated based on PEI attack samples for the hidden encoder $f^{*}$}.$∗ .
If $H^{(i)}_{0}$ is rejected, then $\zeta_{i}$ will be determined to be significantly high and $f_{i}$ will be inferred as the hidden $f^{*}$. We then construct the following statistic z-score to evaluate the null hypothesis,
$z_{i}=\left(\zeta_{i}-\mathrm{E}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}\right)/\left(\mathrm{SD}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}\right),$ (3)
where $\mathrm{E}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}$ and $\mathrm{SD}[\zeta_{i^{\prime}}]_{i^{\prime}=1}^{N}$ are the mean and the sample standard deviation of the $N$ candidate PEI scores $\zeta_{1},\cdots,\zeta_{N}$. If $z_{i}$ is above a chosen threshold, then the null hypothesis $H^{(i)}_{0}$ will be rejected. We further assume that the distribution of PEI scores calculated based on non-hidden encoders’ PEI attack samples can be approximated by a normal distribution. In this case, if we chose $z_{i}>1.7$ as the rejection criterion, it will lead to a one-sided $p$-value, which is also the false-positive rate, of $4.45\%$. Since such a $p$-value is smaller than the typical statistical significance level (which is $5\%$), we thus set the criterion as $z_{i}>1.7$, which eventually leads to the following inference attack,